OpenStudy (anonymous):

I have a question for using quadratic equation again. I have 3^2-5x-7=0 Can any one help me?

OpenStudy (anonymous):

Do you mean 3x^2?

OpenStudy (anonymous):

is nt in the form of quadartic

OpenStudy (anonymous):

yes I meant 3x^2-5x-7=0

OpenStudy (anonymous):

(3x+7)(x-1)

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

x[1] = 1/6 (5-sqrt(109)) x[2] = 1/6 (5+sqrt(109))

OpenStudy (anonymous):

OK doesn't factorise so use formula: x =( 5 +/- sqrt(25 - 4 times 3 times -7))/3 I get x = (5 +/- sqrt109)/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do you know the quadratic formula? The solution of ax^2 + bx + c = 0 is \[x = -b +/- \sqrt{(b^2 - 4ac}) \over 2a\]

OpenStudy (anonymous):

yes I know that

OpenStudy (anonymous):

@ suzi20 would it not be this x=-5+sqrt109/6, x=-5-sqrt109/6

OpenStudy (anonymous):

OK so look at suzi20's correct solution and see if you can get the same answer by substituting for a, b and c in the formula, where a = 3, b = -5 and c = -7

OpenStudy (anonymous):

leahp x=-5+sqrt109/6, x=-5-sqrt109/6 is a correct answer

OpenStudy (anonymous):

thank you, again

OpenStudy (anonymous):

I have one more question does the letter i is not also use for the symbol +- sign together